OpenStudy (anonymous):

Decide if the statement is always, sometimes, or never true. A quadratic equation has two solutions. A. always B. sometimes C. never

OpenStudy (anonymous):

depends on what you mean by \(2\)

OpenStudy (anonymous):

Thats the question i got i don't understand the questions so i came here :(

OpenStudy (anonymous):

counting multiplicity it always has two if you mean two real solutions, then no it could have no real solutions, one real solution or two real solutions

OpenStudy (anonymous):

This is dependent on the value of the discriminant; i.e. the term that appears in the square root of the quadratic formula. If \(ax^2+bx+c\) is our quadratic function, then the discriminant is defined as \(\Delta = b^2-4ac\). Now, we note that if \(\Delta >0\), there's two (unique) solutions, if \(\Delta = 0\), there's one (repeated) solution, and if \(\Delta <0\) there are two imaginary (non real) solutions (i.e. no real solutions).

OpenStudy (anonymous):

nice \(\large \Delta\) !!

OpenStudy (anonymous):

x^2=1, x = -1, +1 x^2 = -1, x = i , -i yes, sometimes real sometimes imaginary

OpenStudy (anonymous):

So that means its B sometimes because depending on the number

OpenStudy (anonymous):

i am willing to be that you math teacher wants "sometimes" wouldn't bet more than $7 though

OpenStudy (anonymous):

Yes; the moral here is that any quadratic can have at most two solutions.

OpenStudy (anonymous):

lol i don't have a teacher my computer is teaching me lol